OpenStudy (anonymous):

What is the 41st term of the arithmetic sequence where a1 = 16 and a15 = -40

hartnn (hartnn):

you know the n'th term formula ?

hartnn (hartnn):

\(\Large a_n = a_1 + (n-1)d\) plug in n = 15, a1 = 1st term = 15 a15 = -40 = left side

OpenStudy (anonymous):

so it would be -40=16+(15-1)d

hartnn (hartnn):

correct! find 'd'

OpenStudy (anonymous):

-4

hartnn (hartnn):

yes, now use the same formula again

hartnn (hartnn):

this time you need 41st term so plug in n=41

hartnn (hartnn):

d=-4 a1 =16

OpenStudy (anonymous):

-144?

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is 14 and the last term is -3,919,104

hartnn (hartnn):

you will need to find 'r' = common ratio first use the formula \(a_n = a_1 r^{n-1}\) n= 8 a1 = 1st term = 14 an = last term = a8 = -3,919,104

OpenStudy (anonymous):

i got like 3700 something weird number

hartnn (hartnn):

for 'r' ?

OpenStudy (anonymous):

yeah lemme try again

hartnn (hartnn):

sure, i am getting an integer a negative integer

OpenStudy (anonymous):

yeah like -3703 ei

hartnn (hartnn):

lol, i am getting r = -6

hartnn (hartnn):

\(-3,919,104 = 14 r^7 \\ \Large r =\sqrt[7]{\dfrac{-3,919,104}{14}} = -6 \)

OpenStudy (anonymous):

oohhhh

OpenStudy (anonymous):

so now don't i plug in -6 into the formula

hartnn (hartnn):

now use the sum formula \(\Large S_n = a_1 \dfrac{r^n-1}{r-1}\) r= -6 a1 =14 n=8

OpenStudy (anonymous):

14 -1679616-1/-7

hartnn (hartnn):

ok, and what does that equal?

OpenStudy (anonymous):

i keep getting 3,359,234

hartnn (hartnn):

\(\huge \checkmark \)

hartnn (hartnn):

thats correct, why do u doubt ?

OpenStudy (anonymous):

cause thats not one of my answers given

hartnn (hartnn):

ohh? let me check all the work again

hartnn (hartnn):

r =- 6 is surely correct

hartnn (hartnn):

for sum i got \(\Large -3359230\)

OpenStudy (anonymous):

okay thats one of them!

OpenStudy (anonymous):

A lottery winner must decide between two methods of payment. Choice one is to receive a lump sum of $100,000,000. Choice two is to begin with one cent on day 1, four cents on day 2, 16 cents on day three, and so on until the end of 16 days. Which choice should the lottery winner select? Using complete sentences, explain the procedure taken to answer this question.

hartnn (hartnn):

can you ask this as a new question ? so that others can help you too ? sorry i need to go now...